[Jenne, Rana, Koji]
Since the MOPA has been having a bad few weeks (and got even more significantly worse in the last day or so), we opened up the MOPA box to increase the power. This involved some adjusting of the NPRO, and some adjusting of the alignment between the NPRO and the Amplifier. Afterward, the power out of the MOPA box was increased. Hooray!
Steps taken:
0. Before we touched anything, the AMPMON was 2.26, PMC_Trans was 2.23, PSL-126MOPA_126MON was 152 (and when the photodiode was blocked, it's dark reading was 23).
1. We took off the side panel of the MOPA box nearest the NPRO, to gain access to the potentiometers that control the NPRO settings. We selectively changed some of the pots while watching PSL-126MOPA_126MON on Striptool.
2. We adjusted the pot labeled "DTEMP" first. (You have to use a dental mirror to see the labels on the PCB, but they're there). We went 3.25 turns clockwise, and got the 126MON to 158.
3. To give us some elbow room, we changed the PSL-126MOPA_126CURADJ from +10.000 to 0.000 so that we have some space to move around on the slider. This changed 126MON to 142. The 126MOPA_CURMON was at 2.308.
4. We tried adjusting the "USR_CUR" pot, which is labeled "POWER" on the back panel of the NPRO (you reach this pot through a hole in the back of the NPRO, not through the side which we took off, like all the other pots today). This pot did nothing at all, so we left it in its original position. This may have been disabled since we use the slider.
5. We adjusted the CUR_SET pot, and got the 126MON up to 185. This changed the 126MOPA_CURMON to 2. 772 and the AMPMON to 2.45
We decided that that was enough fiddling with the NPRO, and moved on to adjusting the alignment into the Amplifier.
6. We teed off of the AMPMON photodiode so that we could see the DC values on a DMM. When we used a T to connect both the DMM and the regular DAQ cable, the DMM read a value a factor of 2 smaller than when the DMM was connected directly to the PD. This shouldn't happen.....it's something on the to-fix-someday list.
7. Rana adjusted the 2 steering mirrors immediately in front of the amplifier, inside the MOPA box. This changed the DMM reading from its original 0.204 to 0.210, and the AMPMON reading from 2.45 to 2.55. While this did help increase the power, the mirrors weren't really moved very much.
8. We then noticed that the beam wasn't really well aligned onto the AMPMON PD. When Rana leaned on the MOPA box, the PD's reading changed. So we moved the PD a little bit to maximize its readings. After this, the AMPMON read 2.68, and the DMM read 0.220.
9. Then Rana adjusted the 2 waveplates in the path from the NPRO to the Amplifier. The first waveplate in the path didn't really change anything. Adjusting the 2nd waveplate gave us an AMPMON of 2.72, and a DMM reading of 0.222.
10. We closed up the MOPA box, and locked the PMC. Unfortunately, the PMC_Trans was only 1.78, down from the 2.26 when we began our activities. Not so great, considering that in the end, the MOPA power went up from 2.26 to 2.72.
11. Koji and I adjusted the steering mirrors in front of the PMC, but we could not get a transmission higher than 1.78.
12. We came back to the control room, and changed the 126MOPA_126CURADJ slider to -2.263 which gives a 126MOPA_CURMON to 2.503. This increased PMC_TRANS up to 2.1.
13. Koji did a bit more steering mirror adjustment, but didn't get any more improvement.
14. Koji then did a scan of the FSS SLOW actuator, and found a better temperature place (~ -5.0)for the laser to sit in. This place (presumably with less mode hopping) lets the PMC_TRANS get up to 2.3, almost 2.4. We leave things at this place, with the 126MOPA_126CURADJ slider at -2.263.
Now that the MOPA is putting out more power, we can adjust the waveplate before the PBS to determine how much power we dump, so that we have ~constant power all the time.
Also, the PMCR view on the Quad TVs in the Control Room has been changed so it actually is PMCR, not PMCT like it has been for a long time. |